OpenStudy (anonymous):

solve for b 0.5(b-30)+0.2b=55

OpenStudy (angela210793):

1/2b-1/2*30+0.2b=55 0.7b=55+15 0.7b=70 b=70/0.7 b=100

OpenStudy (anonymous):

b = 100